please help! no idea where to start on this one
The first apartment I rented was 32 years ago in Chicago. The rent was $250 a
month. If inflation has been more or less around 4% a year and the rent has gone up
according to the inflation. What do you think the rent for that apartment is today?

Answer by Theo(675)

250 a month 32 years ago.

Rent goes up 4% each year.

250 * (1.04)^32 = 877.0146867 today.

Inflation is the increase in price per year.

At 4% per year, the rent is multiplied by 1.04 each year.

first year the rent is $250
Second year the rent is $250 * 1.04 = $260.00
Third year the rent is $260.00 * 1.04 = $270.4

You went up two years.
The formula is $250 * (1.04)^n where n is the number of years.
n = 2
250 * (1.04)^2 = $270.4

When n = 32, the formula becomes:

250 * (1.04)^32 = $877.0146867